# Order isomorphisms of sup-stable function spaces: continuous, Lipschitz, c-convex, and beyond

### v1

@@ Abstract
There have been many parallel streams of research studying order isomorphisms of some specific sets $\mathcal{G}$ of functions from a set $\mathscr{X}$ to {$\mathbb{R}\cup\{\pm\infty\}$, such as the sets of convex or Lipschitz functions. We provide in this article a unified abstract approach inspired by $c$-convex functions. Our results are obtained highlighting the role of inf and sup-irreducible elements of $\mathcal{G}$ and the usefulness of characterizing them, to subsequently derive the structure of order isomorphisms, and in particular of those commuting with the addition of scalars. We show that in many cases all these isomorphisms $J:\mathcal{G}\to\mathcal{G}$ are of the form $Jf=g+f\circ\phi$ for a translation $g:\mathscr{X}\to\mathbb{R}$ and a bijective reparametrization $\phi:\mathscr{X}\to\mathscr{X}$. We apply our theory to the sets of $c$-convex functions on compact Hausdorff spaces, to the set of lower semicontinuous (convex) functions on a Hausdorff topological vector space and to Lipschitz and 1-Lipschitz functions of complete metric spaces.
@@
We are interested in the characterization of the order preserving (anti)-isomorphisms for max or (max,+) operations over some sets $\mathcal{G}$ of functions from a set $\mathscr{X}$ to {$\mathbb{R}\cup\{\pm\infty\}$ such that $\mathcal{G}$ is stable by arbitrary suprema, and possibly by the addition of scalars. Why should one be interested in such characterization of transformations? First because it corresponds to nonlinear versions of the Banach-Stone theorem, and thus relates to studying morphisms between sets, as pursued in [Wea94, LT16]. Second, the characterization shines a new light on the uniqueness of some very familiar transformations such as the Fenchel transform or the polar duality of convex bodies, as done in [AAM09]. Finally, dualities, i.e. order-reversing involutions, are symmetries for some equations like the complex Monge-Ampère [Lem17, BCEKR20] and allow for new Blaschke-Santaló-type inequalities as shown by [AASW23]. The latter two topics are related to the flourishing field of optimal transport, being also targeted by our study with its focus on the set of $c$-convex functions.
@@
Our main result is that, for several function spaces $\mathcal{G}\subset\overline{\mathbb{R}}^{\mathscr{X}}$ of interest, every isomorphism $J$ over $\mathcal{G}$ that commutes with the supremum and the addition of scalars, in short a (max,+)-isomorphism, is of the form (1) owing to the following characterization:
@@
Conversely every $J$ satisfying ii) is a (max,+)-isomorphism.
@@
complete
metric space
Lipschitz functions
Lipschitz bi-Lipschitz homeomorphism 42
@@
continuous affine continuous affine 44
On $c$-convex functions, we show in 27 that the existence of a (max,+)-isomorphism between $b$-convex and $c$-convex functions is equivalent to a “conjugacy” formula (23) satisfied by the kernels. Concerning 1-Lipschitz functions, the results rely on the compactification of metric spaces by horofunctions, and in particular on properties of Busemann functions (36).
Similarly to the proof of the Banach-Stone theorem, see e.g. the lecture notes [GJ02], the key idea in our proofs is to advocate for inf and sup-irreducibility as a good notion of “extremality” (4) and to show that these “extremal” points are preserved by order isomorphisms (13) and generate the whole space, in a Krein-Milman fashion (19 and 25).
Beyond (max,+)-isomorphisms we study the order isomorphisms of sets containing the indicator functions, such as the lower semicontinuous functions on Hausdorff topological spaces (30) and its subset of convex functions (44). In these cases, indicator functions play a major role in the characterization. We in particular obtain an extension of the theorems of [AAM09, IRS15] in the generality of locally convex Hausdorff spaces.
@@
Some of our results have precursors when $\mathscr{X}$ is finite. The notion of irreducibility was used in [BSS07, GK07] when studying the max-plus analogue of the representation of convex sets in terms of extreme points or extreme rays. Concerning infinite sets, the works can be regrouped by the functions spaces studied:
@@
First, on convex functions, our focus meets that of the research program set by S. Artstein-Avidan and V. Milman on the characterization of some operations on subsets of convex functions. In their seminal paper [AAM09], they have shown that, for $\mathcal{G}$ the space of convex lower semicontinuous functions over $\mathbb{R}^{d}$, the only possible anti-involutions are compositions of the Fenchel transform with affine transformations. This result was further extended to Banach spaces by [IRS15], followed by [CL21] who characterized for which Banach spaces the Fenchel transform is an anti-involution. We refer to [IRS15] for a list of other developments on $s$-concave functions or convex bodies. Underlying all of these works on convex functions is a fundamental result of affine geometry that, in dimensions larger than 2, transformation preserving straight lines are affine. The notable difference of our approach with these works is that we tackle spaces beyond convex functions on vector spaces, so in general we do not have straight lines.
@@
Second, we tackle spaces of (lower semi) continuous functions. This is related to developing analogues of the Banach-Stone theorem for order preserving bijections, as for instance done in [KM97, Chapter 2] for the set of continuous functions $C(\mathscr{X},\mathbb{R})$ on a compact set $\mathscr{X}$. The isomorphisms of the set $C(\mathscr{X},\mathbb{R})$ have a long history as nicely recalled on [LT16, p2]. Here we switch instead the focus to the sup-closure of $C(\mathscr{X},\mathbb{R})$, namely the space of lower semicontinuous functions. The isomorphisms of the latter have received less interest so far, and we highlight the role of the indicator functions.
@@
Finally, an important application of ours is to the set of Lipschitz functions on a complete metric space, or its subset of functions with Lipschitz constant at most 1. Most of the interest for spaces of Lipschitz functions has been for their Banach structure, but they also have a rich lattice structure as advocated by [Wea94]. This induces quite some polisemy on the notion of isomorphism. For instance [CCD19] studies Banach space isomorphisms between spaces of Lipschitz functions. [Wea94] leveraged the inf/sup-lattice properties of the vector spaces of Lipschitz functions to characterize linear Banach space isomorphisms preserving Riesz-norms. Banach–Stone-like theorems for Lipschitz functions have also been shown for linear isomorphisms such as [JVCVV09, Theorem 4.1] and references within. Finally Leung et al. for linear, in [LL13], and nonlinear order isomorphisms, in [LT16], designed an abstract framework to describe order isomorphisms on “near vector lattices”. However all the aforementioned authors rely extensively on vector subspaces of the space of continuous functions, while we consider sup-stable sets of functions. The closest research to our work was done by [SS11] for Lipschitz functions on bounded metric spaces and by [CJ17, DJVM20] for $1$-Lipschitz functions on metric spaces with non-symmetric distance with the assumption that the isomorphisms also preserve convex combinations. Nevertheless, none of the authors uses the notions of inf or sup-irreducible functions, whereas these are the backbone of our analysis. We recover many of the results of [LT16] and weaken their assumptions.
@@
In Section 1 we introduce the notions of irreducibility and of isomorphisms that we consider. Section 2 pertains to our main results on order isomorphisms, in particular when commuting with constants, with Section 2.3 devoted to $c$-convex functions. In Section 3, we detail the consequences of these results for continuous and lower semicontinuous functions (Section 3.1), for 1-Lipschitz and Lipschitz functions (Section 3.2), and finally for convex lower semicontinuous functions (Section 3.3).
@@ 1 Preliminaries
Notation: The extended real line is denoted by $\overline{\mathbb{R}}=[-\infty,+\infty]$, we also use the notations $\mathbb{R}_{\top}=(-\infty,+\infty]$ and $\mathbb{R}_{\bot}=[-\infty,+\infty)$. Given any sets $\mathscr{X}$ and $E$, we denote by $E^{\mathscr{X}}$ the set of functions $f$ from $\mathscr{X}$ to $E$. When $\mathscr{X}$ is equipped with a topology, a function $f:\mathscr{X}\rightarrow\overline{\mathbb{R}}$ is said to be lower semicontinuous (l.s.c.) if its epigraph is a closed subset of $\mathscr{X}\times\overline{\mathbb{R}}$. Its domain is defined by $\operatorname*{Dom}(f)=\{x\in\mathscr{X}\,|\,f(x)<\infty\}$. A function $f$ is said to be proper if its domain is non-empty and if it does not take the value $\infty$- ∞. For $\mathcal{G}\subset\overline{\mathbb{R}}^{\mathscr{X}}$ and some $x\in\mathscr{X}$, if there exists a function $g\in\mathcal{G}$ such that $g(x)\in\mathbb{R}$, we say that $\mathcal{G}$ is proper at $x$, and more simply that $\mathcal{G}$ is proper if it is at all $x\in\mathscr{X}$. When $\mathscr{X}$ and $\mathscr{Y}$ are topological vector spaces, the set $\mathscr{X}^{*}$ denotes the dual space of $\mathscr{X}$ of continuous linear forms from $\mathscr{X}$ to $\mathbb{R}$ and $\mathcal{L}(\mathscr{X},\mathscr{Y})$ refers to the space of continuous linear operators from $\mathscr{X}$ to $\mathscr{Y}$. We write the duality product over $\mathscr{X}^{*}\times\mathscr{X}$ as $\langle\cdot,\cdot\rangle:\mathscr{X}^{*}\times\mathscr{X}\to\mathbb{R}$.
@@
We say that a subset $\mathcal{G}$ of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$ is sup-stable if it is stable under arbitrary sups, and that it is a complete subspace of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$ if it is stable under arbitrary sups and addition of scalars. For $\mathcal{G}\subset\overline{\mathbb{R}}^{\mathscr{X}}$, we define its sup-closure $\overline{\mathcal{G}}^{\sup}:=\{\sup_{\alpha\in\mathcal{A}}h_{\alpha}|\,\mathcal{A}\text{ an index set},\,\{h_{\alpha}\}_{\alpha\in\mathcal{A}}\subset\mathcal{G}\}$, and inf-closure $\overline{\mathcal{G}}^{\inf}:=\{\inf_{\alpha\in\mathcal{A}}h_{\alpha}|\,\mathcal{A}\text{ an index set},\,\{h_{\alpha}\}_{\alpha\in\mathcal{A}}\subset\mathcal{G}\}$. If $\mathcal{G}$ is sup-stable, then $\mathcal{G}=\overline{\mathcal{G}}^{\sup}$. We say that $\mathcal{G}$ is inf-stable if $\mathcal{G}=\overline{\mathcal{G}}^{\inf}$.
@@
Let $\mathcal{G}$ be a sup-stable subset of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$. For all $\tilde{g},\tilde{h}\in\overline{\mathbb{R}}^{\mathscr{X}}$, if $-~f=\inf(\tilde{g},\tilde{h})$ with $f\in\mathcal{G}$, then $-~f=\inf\nolimits^{\mathcal{G}}(\inf\nolimits^{\mathcal{G}}\tilde{g},\inf\nolimits^{\mathcal{G}}\tilde{h})$. The $\mathcal{G}$-relatively-inf-irreducible functions are thus also inf-irreducible.
@@
$\delta^{\bot}_{x}\left(y\right):=\left\{\begin{array}[]{ll}0&\text{ if }y=x,\\ -\infty&\text{ otherwise,}\end{array}\right.\delta^{\top}_{x}\left(y\right):=\left\{\begin{array}[]{ll}0&\text{ if }y=x,\\ +\infty&\text{ otherwise.}\end{array}\right.$ (6)
We will often informally refer to them as Dirac masses. Notably, if $\delta^{\top}_{x}\in\mathcal{G}$, then it is clearly $\mathcal{G}$-relatively-inf-irreducible, and if $\delta^{\bot}_{x}\in\mathcal{G}$, then it is sup-irreducible. The term “irreducible” can be understood by analogy with irreducible polynomials.
@@
    a max-isomorphism if it is invertible and if it commutes with suprema, i.e. $J(\sup(f,g))=\sup(Jf,Jg)$, assuming that $\mathcal{G}$ and $\mathscr{F}$ are sup-stable;
@@
Let $\mathscr{F}$ and $\mathcal{G}$ be sup-stable sets. Then the order isomorphisms $J:\mathscr{F}\rightarrow\mathcal{G}$ commute with an arbitrary suprema, and thus coincide with max-isomorphisms. Moreover, if $J:\mathscr{F}\rightarrow\mathcal{G}$ is a max-isomorphism, then so is $J^{-1}:\mathcal{G}\rightarrow\mathscr{F}$.
@@
hence $J(\sup_{\alpha}f_{\alpha})=\sup_{\alpha}Jf_{\alpha}$.
Let $J$ be a max-isomorphism, take $f,g\in\mathcal{G}$, $f_{0},g_{0}\in\mathscr{F}$ such that $f=Jf_{0}$ and $g=Jg_{0}$. Then
@@
The relation between ¯$\overline{\mathbb{R}}_{\max}$-sesquilinear and continuous maps, a.k.a. (Fenchel-Moreau) conjugations, and kernels is one-to-one, so it is equivalent to study an anti-isomorphism or its kernel, as a consequence of
@@
A map $\bar{B}:\overline{\mathbb{R}}^{\mathscr{Y}}\rightarrow\overline{\mathbb{R}}^{\mathscr{X}}$ is ¯$\overline{\mathbb{R}}_{\max}$-sesquilinear and continuous if and only if there exists a kernel $b:\mathscr{X}\times\mathscr{Y}\rightarrow\overline{\mathbb{R}}$ such that $\bar{B}f(x)=\sup_{y\in\mathscr{X}}b(x,y)-f(y)$ for all $f\in\overline{\mathbb{R}}^{\mathscr{Y}}$ (with $\infty$- ∞ absorbing). Moreover in this case $b$ is uniquely determined by $\bar{B}$ as $b(\cdot,y)=\bar{B}\delta^{\top}_{y}$.
9 allows to write
@@
Given a kernel $b:\mathscr{X}\times\mathscr{Y}\to\overline{\mathbb{R}}$, let $\bar{B}:\overline{\mathbb{R}}^{\mathscr{Y}}\rightarrow\overline{\mathbb{R}}^{\mathscr{X}}$ and its transpose $\bar{B}^{\circ}:\overline{\mathbb{R}}^{\mathscr{X}}\rightarrow\overline{\mathbb{R}}^{\mathscr{Y}}$ be defined by
@@
If these properties hold, then $\bar{F}$ can be taken as the restriction of $\bar{B}$ to $\operatorname*{Rg}(B^{\circ})$. Moreover, for $\mathscr{X}=\mathscr{Y}$, there exists an anti-involution $\bar{F}$ over $\mathcal{G}$ if and only if there exists a symmetric $b$ such that $\mathcal{G}=\operatorname*{Rg}(B)$.
This result extends [AFG23, Theorem 5.1] originally written for anti-involutions only. The proof of 10 can be found in the Appendix on pProof of 10.
Anti-involutions were described by [AAM09, Definition 11] as a “concept of duality”, this point of view being developed more at length in [AAM07]. So if there are no anti-involutions over a set $\mathcal{G}$, as that is the case for $\mathcal{G}$ the set of lower semicontinuous functions on a Hausdorff space with no isolated point, as shown in 32, then there is no such duality for $\mathcal{G}$. Describing anti-isomorphisms over complete lattices as an abstract nonlinear duality is actually at the core of abstract convexity theory as summarized in [Sin97]. In this article however, we do not use complete lattices since our sets are not inf-stable, unlike ¯$\overline{\mathbb{R}}^{\mathscr{X}}$, and we focus less on duality and more on order isomorphisms.

### v2

@@ Abstract
There have been many parallel streams of research studying order isomorphisms of some specific sets $\mathcal{G}$ of functions from a set $\mathscr{X}$ to {$\mathbb{R}\cup\{\pm\infty\}$, such as the sets of convex or Lipschitz functions. We propose in this article a unified approach inspired by $c$-convex functions. Our results are obtained highlighting the role of inf and sup-irreducible elements of $\mathcal{G}$ and the usefulness of characterizing them, to subsequently derive the structure of order isomorphisms, and in particular of those commuting with the addition of scalars. We show that in many cases all these isomorphisms $J:\mathcal{G}\to\mathcal{G}$ are of the form $Jf=g+f\circ\phi$ for a translation $g:\mathscr{X}\to\mathbb{R}$ and a bijective reparametrization $\phi:\mathscr{X}\to\mathscr{X}$. Given a reference anti-isomorphism, this characterization then allows to recover all the other anti-isomorphisms. We apply our theory to the sets of $c$-convex functions on compact Hausdorff spaces, to the set of lower semicontinuous (convex) functions on a Hausdorff topological vector space and to 1-Lipschitz functions of complete metric spaces. The latter application is obtained using properties of the horoboundary of a metric space.
Keywords: order isomorphisms, 1-Lipschitz functions, convex functions.
MSC classes: 15A80, 06D50, 26B25, 06A15.
@@
We are interested in the characterization of the order preserving (anti)-isomorphisms for max or (max,+) operations over some sets $\mathcal{G}$ of functions from a set $\mathscr{X}$ to {$\mathbb{R}\cup\{\pm\infty\}$ such that $\mathcal{G}$ is stable by arbitrary suprema, and possibly by the addition of scalars. Why should one be interested in such characterization of transformations? First because it corresponds to nonlinear versions of the Banach-Stone theorem, and thus relates to studying morphisms between sets, as pursued in [Wea94, LT16]. Second, the characterization shines a new light on the uniqueness of some very familiar transformations such as the Fenchel transform [AAM09] or the polar duality of convex bodies [AAM08, BS08], especially in the context of valuation theory in convex geometry [Lud23]. Finally, dualities, i.e. order reversing involutions, are symmetries for some equations like the complex Monge-Ampère [Lem17, BCEKR20] and allow for new Blaschke-Santaló-type inequalities as shown by [AASW23]. The latter two topics are related to the flourishing field of optimal transport, being also targeted by our study with its focus on the set of $c$-convex functions.
@@
We show that, for several function spaces $\mathcal{G}\subset\overline{\mathbb{R}}^{\mathscr{X}}$ of interest, every bijection $J$ over $\mathcal{G}$ that commutes with the pairwise supremum and the addition of scalars, in short a (max,+)-isomorphism, is of the form (1) owing to the following characterization:
@@
Conversely every $J$ satisfying ii) is a (max,+)-isomorphism, and, the representation $(g,\phi)$ in (1) is unique.
@@
continuous affine
invertible affine
with w*-w* adjoint
42
Concerning order reversing isomorphisms $T$ between sets $\mathscr{F}$ and $\mathcal{G}$ commuting with the addition of scalars, we characterize their existence through kernels $b$ in 10. Owing to the immediate 11, the uniqueness of $T$ up to basic transformations is instead a consequence of results like 1.
On $c$-convex functions over a compact set, we show in 27 that the existence of a (max,+)-isomorphism between $b$-convex and $c$-convex functions is equivalent to a “conjugacy” formula (25) satisfied by the kernels. Concerning 1-Lipschitz functions, the results rely on the compactification of metric spaces by horofunctions, and in particular on properties of Busemann functions (36), we also allow for weak distances, i.e. potentially assymetric and with negative values.
Similarly to the proof of the Banach-Stone theorem, see e.g. the lecture notes [GJ02], the key idea in our proofs is to advocate for inf and sup-irreducibility as a good notion of “extremality” (4) and to show that these “extremal” points are preserved by order isomorphisms (13) and generate the whole space, in a Krein–Milman fashion (19 and 25).
Beyond (max,+)-isomorphisms, we study the order isomorphisms of sets containing the indicator functions of points, such as the set of lower semicontinuous functions on Hausdorff topological spaces (30) and its subset of convex functions (42). In these cases, indicator functions play a major role in the characterization. We in particular obtain an extension of the theorems of [AAM09, IRS15] in the generality of locally convex Hausdorff spaces.
@@
Some of our results have precursors when $\mathscr{X}$ is finite, in the setting of tropical geometry – the (max,+) structure considered here being the main example of tropical algebraic structure. In particular, the notion of irreducibility was used in [BSS07, GK07] when studying the max-plus analogue of the representation of convex sets in terms of extreme points or extreme rays. Concerning infinite sets, the literature can be regrouped according to the functions spaces studied:
@@
First, on convex functions, our focus meets that of the research program set by S. Artstein-Avidan and V. Milman on the characterization of some operations on subsets of convex functions. In their seminal paper [AAM09], they have shown that, for $\mathcal{G}$ the space of convex lower semicontinuous functions over $\mathbb{R}^{d}$, the only possible anti-involutions are compositions of the Fenchel transform with affine transformations. This result was further extended to Banach spaces by [IRS15], followed by [CL21] who characterized for which Banach spaces the Fenchel transform is an anti-involution. We refer to [IRS15, AASW23] for a list of other developments on $s$-concave functions or convex bodies. Underlying all of these works on convex functions is a fundamental result of affine geometry that, in dimensions larger than 2, transformations preserving straight lines are affine. The notable difference of our approach is that we consider spaces beyond convex functions on vector spaces, so in general we do not have linear structures.
@@
Second, we tackle spaces of (lower semi) continuous functions. This is related to developing analogues of the Banach-Stone theorem for order preserving bijections, as for instance done for the set of continuous functions $C(\mathscr{X},\mathbb{R})$ on a compact set $\mathscr{X}$ in [Kap48, CS08] or in the max-plus literature [KM97, Chapter 2]. The isomorphisms of the set $C(\mathscr{X},\mathbb{R})$ and its subsets have indeed a long history based on many viewpoints, as nicely recalled in [LT16, p2]. Some interpretations can be more insightful, for instance, seeing $C(\mathscr{X},\mathbb{R})$ as a vector lattice, Kaplansky’s Theorem gives the famed Gelfand-Kolmogorov theorem which focuses on $C(\mathscr{X},\mathbb{R})$ as an algebra. Here we switch instead the focus to the sup-closure of $C(\mathscr{X},\mathbb{R})$, namely the space of lower semicontinuous functions. The isomorphisms of the latter have received less interest so far, and we highlight the role of the indicator functions.
@@
Finally, an important application of ours is to the set of Lipschitz functions on a complete metric space with Lipschitz constant at most 1. Most of the interest for spaces of Lipschitz functions has been for their Banach structure, but they also have a rich lattice structure as advocated by [Wea94]. This induces quite some polisemy on the notion of isomorphism. For instance [CCD19] studies Banach space isomorphisms between spaces of Lipschitz functions. Weaver leveraged instead the inf/sup-lattice properties of the vector spaces of Lipschitz functions to characterize linear Banach space isomorphisms preserving Riesz-norms [Wea94]. Banach–Stone-like theorems for Lipschitz functions have also been shown for linear isomorphisms such as [JVCVV09, Theorem 4.1] and nonlinear [CSCS11, Theorem 1] and references within. Finally Leung et al. for linear, in [LL13], and nonlinear order isomorphisms, in [LT16], designed an abstract framework to describe order isomorphisms on “near vector lattices”. However all the aforementioned authors rely extensively on vector subspaces of the space of continuous functions, while we consider sup-stable sets of functions. The closest research to our work was done by [CSJ17, DJVM20] for $1$-Lipschitz functions on metric spaces with non-symmetric distance with the assumption that the isomorphisms also preserve convex combinations. Nevertheless, none of the authors uses the notions of inf or sup-irreducible functions, whereas these are the backbone of our analysis.
@@
In Section 1 we introduce the notions of irreducibility and of isomorphisms that we consider. Section 2 pertains to our main results on order isomorphisms, in particular when commuting with constants, with Section 2.3 devoted to $c$-convex functions. In Section 3, we detail the consequences of these results for continuous and lower semicontinuous functions (Section 3.1), for 1-Lipschitz functions (Section 3.2), and finally for convex lower semicontinuous functions (Section 3.3).
@@ 1 Preliminaries
Notation: The extended real line is denoted by $\overline{\mathbb{R}}=[-\infty,+\infty]$, we also use the notations $\mathbb{R}_{\top}=(-\infty,+\infty]$, $\mathbb{R}_{\bot}=[-\infty,+\infty)$ and $\mathbb{R}_{>0}=(0,+\infty)$. Given any sets $\mathscr{X}$ and $E$, we denote by $E^{\mathscr{X}}$ the set of functions $f$ from $\mathscr{X}$ to $E$ and $\operatorname*{Id}_{\mathscr{X}}:\mathscr{X}\to\mathscr{X}$ the identity map. When $\mathscr{X}$ is equipped with a topology, a function $f:\mathscr{X}\rightarrow\overline{\mathbb{R}}$ is said to be lower semicontinuous (l.s.c.) if its epigraph is a closed subset of $\mathscr{X}\times\overline{\mathbb{R}}$. Its domain is defined by $\operatorname*{Dom}(f)=\{x\in\mathscr{X}\,|\,f(x)<+\infty\}$. A function $f$ is said to be proper if its domain is non-empty and if it does not take the value $\infty$- ∞. For $\mathcal{G}\subset\overline{\mathbb{R}}^{\mathscr{X}}$ and some $x\in\mathscr{X}$, if there exists a function $g\in\mathcal{G}$ such that $g(x)\in\mathbb{R}$, we say that $\mathcal{G}$ is proper at $x$, and more simply that $\mathcal{G}$ is proper if it is at all $x\in\mathscr{X}$. When $\mathscr{X}$ and $\mathscr{Y}$ are topological vector spaces, the set $\mathscr{X}^{*}$ denotes the dual space of $\mathscr{X}$ of continuous linear forms from $\mathscr{X}$ to $\mathbb{R}$ and $\mathcal{L}(\mathscr{X},\mathscr{Y})$ refers to the space of continuous linear operators from $\mathscr{X}$ to $\mathscr{Y}$. We write the duality product over $\mathscr{X}^{*}\times\mathscr{X}$ as $\langle\cdot,\cdot\rangle:\mathscr{X}^{*}\times\mathscr{X}\to\mathbb{R}$.
@@
We say that a subset $\mathcal{G}$ of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$ is sup-stable (resp. finitely sup-stable) if it is stable under arbitrary sups (resp. pairwise sups), and that it is a complete subspace of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$ if it is stable under arbitrary sups and addition of scalars. For $\mathcal{G}\subset\overline{\mathbb{R}}^{\mathscr{X}}$, we define its sup and inf-closures as
$\overline{\mathcal{G}}^{\sup}$ :=$:=\{\sup_{\alpha\in\mathcal{A}}h_{\alpha}\,|\,\mathcal{A}\text{ an index set},\,\{h_{\alpha}\}_{\alpha\in\mathcal{A}}\subset\mathcal{G}\}$
$\overline{\mathcal{G}}^{\inf}$ :=$:=\{\inf_{\alpha\in\mathcal{A}}h_{\alpha}\,|\,\mathcal{A}\text{ an index set},\,\{h_{\alpha}\}_{\alpha\in\mathcal{A}}\subset\mathcal{G}\}$
The set $\mathcal{G}$ is sup-stable if and only if $\mathcal{G}=\overline{\mathcal{G}}^{\sup}$. We say that $\mathcal{G}$ is inf-stable if $\mathcal{G}=\overline{\mathcal{G}}^{\inf}$.
@@
The term “irreducible” can be understood by analogy with irreducible polynomials.
@@
Let $\mathcal{G}$ be a sup-stable subset of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$. For all $\tilde{g},\tilde{h}\in\overline{\mathbb{R}}^{\mathscr{X}}$, if $f=\inf(\tilde{g},\tilde{h})$ with $f\in\mathcal{G}$, then $f=\inf\nolimits^{\mathcal{G}}(\inf\nolimits^{\mathcal{G}}\tilde{g},\inf\nolimits^{\mathcal{G}}\tilde{h})$. The $\mathcal{G}$-relatively-inf-irreducible functions are thus also inf-irreducible.
@@
$\delta^{\bot}_{x}\left(y\right):=\left\{\begin{array}[]{ll}0&\text{ if }y=x,\\ -\infty&\text{ otherwise,}\end{array}\right.\quad\delta^{\top}_{x}\left(y\right):=\left\{\begin{array}[]{ll}0&\text{ if }y=x,\\ +\infty&\text{ otherwise.}\end{array}\right.$ (6)
We will often informally refer to them as Dirac masses. Notably, if $\delta^{\top}_{x}\in\mathcal{G}$, then it is clearly $\mathcal{G}$-relatively-inf-irreducible, and if $\delta^{\bot}_{x}\in\mathcal{G}$, then it is sup-irreducible.
@@
    a max-isomorphism if it is invertible and if it commutes with pairwise suprema, i.e. $J(\sup(f,g))=\sup(Jf,Jg)$, assuming that $\mathcal{G}$ and $\mathscr{F}$ are sup-stable;
@@
Let ¯$\mathscr{F}\subset\overline{\mathbb{R}}^{\mathscr{Y}}$ and $\mathcal{G}\subset\overline{\mathbb{R}}^{\mathscr{X}}$ be sup-stable sets. Then the order isomorphisms $J:\mathscr{F}\rightarrow\mathcal{G}$ commute with an arbitrary suprema, and thus coincide with max-isomorphisms. Moreover, if $J:\mathscr{F}\rightarrow\mathcal{G}$ is a max-isomorphism, then so is $J^{-1}:\mathcal{G}\rightarrow\mathscr{F}$. Anti-isomorphisms $\bar{F}:\mathscr{F}\to\mathcal{G}$ send relative infima of $\mathscr{F}$ onto suprema of $\mathcal{G}$, i.e. $\bar{F}(\inf\nolimits^{\mathscr{F}}_{\alpha\in\mathscr{A}}f_{\alpha})=\sup_{\alpha\in\mathscr{A}}\bar{F}(f_{\alpha})$ for any $\{f_{\alpha}\}_{\alpha\in\mathscr{A}}\subset\mathscr{F}$.
@@
hence $J(\sup_{\alpha}f_{\alpha})=\sup_{\alpha}Jf_{\alpha}$. Let $J$ be a max-isomorphism, take $f,g\in\mathcal{G}$, $f_{0},g_{0}\in\mathscr{F}$ such that $f=Jf_{0}$ and $g=Jg_{0}$. Then
@@
Similarly, let $\bar{F}^{-1}:\mathcal{G}\to\mathscr{F}$ be the inverse of $\bar{F}$. As $\bar{F}$ is an order anti-isomorphism, it satisfies that, as $\inf\nolimits^{\mathscr{F}}_{\alpha\in\mathscr{A}}f_{\alpha}\leq f_{\beta}$ for any $\beta\in\mathscr{A}$,
Since $\mathcal{G}$ is sup-stable, $\sup_{\alpha\in\mathscr{A}}\bar{F}(f_{\alpha})\in\mathcal{G}$, so composing by $\bar{F}^{-1}$ and using that $\bar{F}^{-1}$ is order reversing, we derive that
$\bar{F}^{-1}\bar{F}(\inf\nolimits^{\mathscr{F}}_{\alpha\in\mathscr{A}}f_{\alpha})\leq\bar{F}^{-1}(\sup_{\alpha\in\mathscr{A}}\bar{F}(f_{\alpha}))\leq\bar{F}^{-1}\bar{F}(f_{\beta}).$
$\inf\nolimits^{\mathscr{F}}_{\alpha\in\mathscr{A}}f_{\alpha}\leq\bar{F}^{-1}(\sup_{\alpha\in\mathscr{A}}\bar{F}(f_{\alpha}))\leq\inf\nolimits^{\mathscr{F}}_{\beta\in\mathscr{A}}f_{\beta}.$
So $\mathscr{F}(\inf\nolimits^{\mathscr{F}}_{\alpha\in\mathscr{A}}f_{\alpha})=\sup_{\alpha\in\mathscr{A}}\bar{F}(f_{\alpha})$.
@@
A map $\bar{B}:\overline{\mathbb{R}}^{\mathscr{Y}}\rightarrow\overline{\mathbb{R}}^{\mathscr{X}}$ is ¯$\overline{\mathbb{R}}_{\max}$-sesquilinear and continuous if and only if there exists a kernel $b:\mathscr{X}\times\mathscr{Y}\rightarrow\overline{\mathbb{R}}$ such that $\bar{B}f(x)=\sup_{y\in\mathscr{Y}}b(x,y)-f(y)$ for all $f\in\overline{\mathbb{R}}^{\mathscr{Y}}$ (with $\infty$- ∞ absorbing). Moreover in this case $b$ is uniquely determined by $\bar{B}$ as $b(\cdot,y)=\bar{B}\delta^{\top}_{y}$.
Hence the relation between ¯$\overline{\mathbb{R}}_{\max}$-sesquilinear and continuous maps, a.k.a. (Fenchel-Moreau) conjugations, and kernels is one-to-one, so it is equivalent to study a map or its kernel. Moreover 9 allows to write
@@
Obviously there may exist some more parcimonious representatives through other $b$ and $\mathscr{Y}$. Given a kernel $b:\mathscr{X}\times\mathscr{Y}\to\overline{\mathbb{R}}$, let $\bar{B}:\overline{\mathbb{R}}^{\mathscr{Y}}\rightarrow\overline{\mathbb{R}}^{\mathscr{X}}$ and its transpose $\bar{B}^{\circ}:\overline{\mathbb{R}}^{\mathscr{X}}\rightarrow\overline{\mathbb{R}}^{\mathscr{Y}}$ be defined by
@@
If these properties hold, then $\bar{F}$ can be taken as the restriction of $\bar{B}:\overline{\mathbb{R}}^{\mathscr{Y}}\to\overline{\mathbb{R}}^{\mathscr{X}}$ to $\operatorname*{Rg}(B^{\circ})$. Moreover, for $\mathscr{X}=\mathscr{Y}$, there exists an anti-involution $\bar{F}$ over $\mathcal{G}$ if and only if there exists a symmetric $b$ such that $\mathcal{G}=\operatorname*{Rg}(B)$.
This result extends [AFG24, Theorem 5.1] originally written for anti-involutions only. The proof of 10 can be found in the Appendix on pProof of 10. 10 exclusively provides the existence of the anti-isomorphism and does not discuss its uniqueness, which in general does not hold. It is one of the goals of this article to show that in many cases the map $\bar{F}$ is nevertheless unique up to a translation $g$ and reparametrization $\phi$. Anti-involutions were described by [AAM09, Definition 11] as a “concept of duality”, this point of view being developed more at length in [AAM07]. 10 can also be seen as a functional analogue to [AASW23, Theorem 1.6] written for the power set $2^{\mathscr{X}}$. If there are no anti-involutions over a set $\mathcal{G}$, as that is the case for $\mathcal{G}$ the set of lower semicontinuous functions on a Hausdorff space with no isolated point, as shown in 32, then there is no such duality for $\mathcal{G}$. Describing anti-isomorphisms over complete lattices as an abstract nonlinear duality is actually at the core of abstract convexity theory as summarized in [Sin97]. In this article however, we do not use complete lattices since our sets are not inf-stable, unlike ¯$\overline{\mathbb{R}}^{\mathscr{X}}$, and we focus less on dualities and more on order isomorphisms.